# MATTER: Multiscale Attention for Registration Error Regression

### v1

@@ 1 Introduction
Point cloud registration (PCR) is a fundamental task in three-dimensional (3D) computer vision, aiming to estimate the rigid transform between two overlapping point clouds. It serves as a backbone for numerous downstream applications, including simultaneous localization and mapping (SLAM), 3D reconstruction, and robotic navigation. While recent advances in PCR have led to significant improvements in accuracy and robustness, registration error is still significant in challenging scenarios, e.g., when optimization of the loss function gets trapped in local minima [1], under motion distortion [2] or measurement noise, or when the problem is geometrically under-constrained [3, 4]. Errors in PCR can propagate to downstream modules, resulting in map distortion or navigation failure. This motivates the need for registration misalignment detection, i.e., predicting, after registration, whether the registered point clouds deviate from the well-aligned ground truth. This prediction can then allow downstream tasks to take corrective action.
@@ 2.3 Feature Extraction
For each pair $(P^{\mathrm{S}},P^{\mathrm{R}})$, we first transform both point clouds into a common frame and use farthest point sampling [14] to obtain a set of $N$ anchor points from $P^{\mathrm{S}}\cup P^{\mathrm{R}}$. In the rest of the section, we focus on a generic anchor point. Let $\mathcal{N}_{P^{\mathrm{S}}}$ and $\mathcal{N}_{P^{\mathrm{R}}}$ denote the points from the source and reference point clouds, respectively, within the neighborhood of this anchor point. We use $\mathcal{N}_{\mathrm{sep}}$ to refer to $\mathcal{N}_{P^{\mathrm{S}}}$ if the anchor point belong to $P^{\mathrm{S}}$, and to $\mathcal{N}_{P^{\mathrm{R}}}$ if it belongs to $P^{\mathrm{R}}$. Define the joint neighborhood $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\mathrm{S}}}\cup\mathcal{N}_{P^{\mathrm{R}}}$. Following [11], we extract local features around each anchor, namely the separate and joint differential entropy [13] $H(\mathcal{N}{\mathrm{sep}}),H(\mathcal{N}{\mathrm{joint}})$, and the Sinkhorn divergence $D_{\lambda}$. They measure the discrepancy between the distributions of points in the aligned source with respect to reference neighborhoods. Additional details on the feature computation are provided in Appendix A. Unreliable neighborhoods are handled explicitly: if one point cloud contributes no points, the differential entropy and Sinkhorn divergence are set to a bounded default. The next local features are coverage ratios $\rho_{\mathrm{joint}}=\big{|}\mathcal{N}_{\mathrm{joint}}\big{|}/(\big{|}P^{\mathrm{S}}\big{|}+\big{|}P^{\mathrm{R}}\big{|})$ and $\rho_{\mathrm{sep}}=\big{|}\mathcal{N}_{\mathrm{sep}}\big{|}/\big{|}P_{\mathrm{sep}}\big{|}$, where $\big{|}P_{\mathrm{sep}}\big{|}$ is given by $\big{|}P^{\mathrm{{S}}}\big{|}$ or $\big{|}P^{\mathrm{{R}}}\big{|}$ if the anchor point belongs to $P^{\mathrm{{S}}}$ or $P^{\mathrm{{R}}}$, respectively.
@@ 2.4 Multiscale Cross-Attention
Multiscale Features. In FACT [11], the radius of $\mathcal{N}_{\mathrm{joint}}$ is fixed and empirically optimized. However, there is a tradeoff between using small neighborhoods and large ones. A small radius may fail when the initial alignment is poor, since the corresponding structure can fall outside the neighborhoods; in contrast, a large neighborhood contains non-overlapping regions, which makes the entropy and Sinkhorn divergence less reliable for estimating registration errors. (We will illustrate this in Fig. 2.) Moreover, a larger neighborhood increases distributional heterogeneity and thus deviates from the approximately isotropic Gaussian assumption. This undermines the validity of the closed-form differential-entropy estimate.
To resolve these failure modes, we compute features at $S$ radii for every anchor point. For each scale $s\in\{1,\dots,S\}$, we query the source and reference clouds to form separate neighborhoods $\mathcal{N}_{P^{\rm S}}^{(s)}$, $\mathcal{N}_{P^{\rm R}}^{(s)}$ and their union $\mathcal{N}_{\text{joint}}^{(s)}=\mathcal{N}_{P^{\rm S}}^{(s)}\cup\mathcal{N}_{P^{\rm R}}^{(s)}$. On these sets we compute the same family of features as in the single scale case. The overall feature vector for this anchor point is denoted by $\mathbf{f}\in\mathbb{R}^{5S+3}$. It contains five per-scale features $\{H(\mathcal{N}_{\mathrm{joint}}^{(s)}),H(\mathcal{N}_{\mathrm{sep}}^{(s)}),D_{\lambda}^{(s)},\rho_{\mathrm{joint}}^{(s)},\rho_{\mathrm{sep}}^{(s)}\}$, and three global features $\{c,d,b\}$.
Multiscale Attention. To further alleviate the trade-off of neighborhood scale, we use multiscale attention learned from the feature vector to weight the scale-dependent features. Specifically, we split the $(5S+3)$-dimensional feature vector $\mathbf{f}$ into three 8-dimensional parts, each containing 5 scale-specific features of a given scale and the 3 global features:
@@ 3.2 Evaluation Results
We train and evaluate on the three datasets described above and report the root mean square error (RMSE), the mean absolute error (MAE), and the coefficient of determination ($R^{2}$). For our method, we consider three radii $r_{1}=7.5$ $\text{m},$ $r_{2}=4.0\,\text{m},$ $r_{3}=2.5\,\text{m}$, and use temperature $\tau=0.6$. We report the results of our method with different $\tau$ in Appendix B.
For all baselines, we follow the same settings as reported in the original papers. E.g. FACT and CorAl both use the fixed radius $7.5\,\mathrm{m}$. Quantitative results in Table 1 show that our method achieves the best performance on all datasets in terms of all three metrics RMSE, MAE, and $R^{2}$. Notably, on nuScenes-ICP, most methods have $\mathrm{RMSE}$ slightly above $\mathrm{MAE}$, indicating few large outliers. In contrast, on KITTI–GeoTransformer, $\mathrm{RMSE}$ is markedly larger than $\mathrm{MAE}$, indicating heavier-tailed errors that are harder to detect.
To better understand the cause of the improvement, we visualize in Fig. 2 the registered point clouds and the per-anchor-point scale that is most highly weighted by attention. The mechanism favors the smallest scale S1 in dense, well-overlapping regions where the local geometry better approximates a 3D Gaussian, making the closed-form differential-entropy descriptors more reliable. When correspondences are absent or alignment is poor, the model adaptively enlarges the radius. In Fig. 2(2(a)), the left side of the reference point cloud lacks corresponding points from the source, which triggers the largest scale S3 so that anchors still cover areas that contain matches. In Fig. 2(2(b)), the largest scale S3 is favored in the red circles because small neighborhoods fail to capture counterparts of both clouds. The midscale S2 is generally less preferred but appears in cluttered medium-thickness structures. In short, the point-wise multiscale selection uses small neighborhoods under high overlap and expands them under low overlap, outperforming single-scale designs.
@@
(a) Non-overlapping region on the bottom-left. Colored circles visualize the neighborhood ranges of the corresponding scales.
(b) Case for poor registration with error = $2.46$ m The red circle zooms in a challenging region where larger neighborhoods are preferred.
@@
We report an ablation on the radius of the neighborhood and multiscale attention, and summarize the results in Table 2. Single-scale models prefer different radii per dataset: 7.5 m in nuScenes (high overlap) and 2.5 m in KITTI (low/moderate overlap), revealing the limitation of a fixed scale. Plain multiscale concatenation already improves over single-scale on nuScenes (noisy) and KITTI, but our point-wise scale-selector attention achieves the lowest RMSE on all datasets with only approximately 0.06% more parameters.
@@ Appendix A Feature Extraction
For each pair $(P^{\mathrm{S}},P^{\mathrm{R}})$, we first transform both point clouds into a common frame and use farthest point sampling [14] to obtain a subset of anchor points.
Differential Entropy. Following Adolfsson et al. [13], we use differential entropy as features. Let $\mathcal{N}_{P^{\rm S}}$ and $\mathcal{N}_{P^{\rm R}}$ denote the points from the source and reference scans within the neighborhood of anchor point $p$, and let the joint neighborhood be $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\rm S}}\cup\mathcal{N}_{P^{\rm R}}$. Assuming that each neighborhood has an approximately isotropic Gaussian distribution, its differential entropy has a closed form.
@@
where $\Sigma$ is the sample covariance of the points in neighborhood $\mathcal{N}$. We compute $H(\mathcal{N}_{\mathrm{sep}})$ for the anchor’s neighborhood in its native cloud ($\mathcal{N}_{\mathrm{sep}}=\mathcal{N}_{P^{\rm S}}$ or $\mathcal{N}_{P^{\rm R}}$, depending on which point cloud the anchor point belongs to), and the joint entropy $H(\mathcal{N}_{\mathrm{joint}})$. Well-aligned scans yield small gaps between $H(\mathcal{N}_{\mathrm{joint}})$ and $H(\mathcal{N}_{\mathrm{sep}})$, whereas misalignment inflates dispersion when the two sets are merged, increasing both gaps.
@@
$\quad-\tfrac{1}{2}\Big{[}W_{\lambda}(\mathcal{N}_{P^{S}},\mathcal{N}_{P^{S}})+W_{\lambda}(\mathcal{N}_{P^{R}},\mathcal{N}_{P^{R}})\Big{]}~,$ (9)

### v2

@@ 1 Introduction
Point cloud registration (PCR) is a fundamental task in three-dimensional (3D) computer vision, aiming to estimate the rigid transform between two overlapping point clouds. It serves as a backbone for numerous downstream applications, including simultaneous localization and mapping (SLAM), 3D reconstruction, and robotic navigation. Despite recent advances in PCR, registration error is still significant in challenging scenarios, e.g., when optimization of the loss function gets trapped in local minima [1], under motion distortion [2] or measurement noise, or when the problem is geometrically under-constrained [3, 4]. Errors in PCR can propagate to downstream modules, resulting in map distortion and navigation failure. This motivates the need for registration misalignment detection, i.e., predicting, after registration, whether the registered point clouds deviate from the well-aligned ground truth. This prediction can then allow downstream tasks to take corrective action.
@@ 2.3 Feature Extraction
We first transform both $P^{\mathrm{S}}$ and $P^{\mathrm{R}}$ point clouds into a common frame and apply the farthest point sampling [14] to select the same number of anchor points from each of them. In the rest of the section, we focus on a generic anchor point. Let $\mathcal{N}_{P^{\mathrm{S}}}$ and $\mathcal{N}_{P^{\mathrm{R}}}$ denote the points from the source and reference point clouds, respectively, within the neighborhood of this anchor point. We use $\mathcal{N}_{\mathrm{sep}}$ to refer to $\mathcal{N}_{P^{\mathrm{S}}}$ if the anchor point belong to $P^{\mathrm{S}}$, and to $\mathcal{N}_{P^{\mathrm{R}}}$ if it belongs to $P^{\mathrm{R}}$. Define the joint neighborhood $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\mathrm{S}}}\cup\mathcal{N}_{P^{\mathrm{R}}}$. Following [11], we extract local features around each anchor, namely the separate and joint differential entropy [13] $H(\mathcal{N}{\mathrm{sep}}),H(\mathcal{N}{\mathrm{joint}})$, and the Sinkhorn divergence $D_{\lambda}$. They measure the discrepancy between the distributions of points in the aligned source with respect to reference neighborhoods. Additional details on the feature computation are provided in Appendix. A. Unreliable neighborhoods are handled explicitly: if one point cloud contributes no points, the differential entropy and Sinkhorn divergence are set to a bounded default. The next local features are coverage ratios $\rho_{\mathrm{joint}}=\big|\mathcal{N}_{\mathrm{joint}}\big|/(\big|P^{\mathrm{S}}\big|+\big|P^{\mathrm{R}}\big|)$ and $\rho_{\mathrm{sep}}=\big|\mathcal{N}_{\mathrm{sep}}\big|/\big|P_{\mathrm{sep}}\big|$, where $\big|P_{\mathrm{sep}}\big|$ is given by $\big|P^{\mathrm{{S}}}\big|$ or $\big|P^{\mathrm{{R}}}\big|$ if the anchor point belongs to $P^{\mathrm{{S}}}$ or $P^{\mathrm{{R}}}$, respectively.
@@ 2.4 Multiscale Cross-Attention
Multiscale Features. In FACT [11], the radius of $\mathcal{N}_{\mathrm{joint}}$ is chosen as a single value, adaptively adjusted according to $d$. There is a tradeoff between using small neighborhoods and large ones. A small radius may fail when the initial alignment is poor, since the corresponding structure can fall outside the neighborhoods; in contrast, a large neighborhood contains non-overlapping regions, which makes the entropy and Sinkhorn divergence less reliable for estimating registration errors. (We will illustrate this in Fig. 2.) Adapting a single radius according to $d$, as in FACT [11], does not capture well this tradeoff. Moreover, a larger neighborhood increases distributional heterogeneity and thus deviates from the approximately isotropic Gaussian assumption. This undermines the validity of the closed-form differential-entropy estimate.
To resolve these failure modes, we compute features at $S$ different radii for every anchor point, and fuse these features with attention. For each scale $s\in\{1,\dots,S\}$, we query the source and reference clouds to form separate neighborhoods $\mathcal{N}_{P^{\rm S}}^{(s)}$, $\mathcal{N}_{P^{\rm R}}^{(s)}$ and their union $\mathcal{N}_{\text{joint}}^{(s)}=\mathcal{N}_{P^{\rm S}}^{(s)}\cup\mathcal{N}_{P^{\rm R}}^{(s)}$. On these sets we compute the same family of features as in the single scale case. The overall feature vector for this anchor point is denoted by $\mathbf{f}\in\mathbb{R}^{5S+3}$. It contains five per-scale features $\{H(\mathcal{N}_{\mathrm{joint}}^{(s)}),H(\mathcal{N}_{\mathrm{sep}}^{(s)}),D_{\lambda}^{(s)},\rho_{\mathrm{joint}}^{(s)},\rho_{\mathrm{sep}}^{(s)}\}$, and three global features $\{c,d,b\}$.
Multiscale Attention. To further alleviate the trade-off of neighborhood scale, we use multiscale attention learned from the feature vector to weight the scale-dependent features. Specifically, we split the $(5S+3)$-dimensional feature vector $\mathbf{f}$ into $S$ 8-dimensional parts, each containing 5 scale-specific features of a given scale and the 3 global features:
@@ 3.2 Evaluation Results
We train and evaluate on the three datasets described above and report the root mean square error (RMSE), the mean absolute error (MAE), and the coefficient of determination ($R^{2}$). For our method, we consider three radii $r_{1}=7.5$ $\mathrm{m},$ $r_{2}=4.0\,\mathrm{m},$ $r_{3}=2.5\,\mathrm{m}$, and use temperature $\tau=0.6$. We report the results of our method with different $\tau$ in Appendix. B.
For all baselines, we follow the same settings as reported in the original papers, e.g., FACT and CorAl both use a single adaptive radius within $[0.5,7.5]$ m. Results in Table 1 show that our method achieves the best performance on all datasets in terms of all three metrics RMSE, MAE, and $R^{2}$. Notably, on nuScenes-ICP, most methods have $\mathrm{RMSE}$ slightly above $\mathrm{MAE}$, indicating few large outliers. In contrast, on KITTI–GeoTransformer, $\mathrm{RMSE}$ is markedly larger than $\mathrm{MAE}$, indicating heavier-tailed errors that are harder to detect.
To better understand the cause of the improvement, we visualize the the registered point clouds in Fig. 2, and the per-anchor-point scale that is most highly weighted by MATTER’s attention. MATTER adaptively favors different radii, while FACT sets its single radius to $7.5$ m for all anchor points in Fig. 2(2(a)) and $95\%$ of the anchor points in Fig. 2(2(b)). Specifically, our attention mechanism favors the smallest scale S1 in dense, well-overlapping regions where the local geometry better approximates a 3D Gaussian, making the closed-form differential-entropy descriptors more reliable. When correspondences are absent or alignment is poor, the model adaptively enlarges the radius. In Fig. 2(2(a)), the left side of the reference point cloud lacks corresponding points from the source, which triggers the largest scale S3 so that anchors still cover areas that contain matches. In Fig. 2(2(b)), the largest scale S3 is favored in the red circles because small neighborhoods fail to capture counterparts of both clouds. The midscale S2 is generally less preferred but appears in cluttered medium-thickness structures.
@@
(a) The case of poor registration with $E_{\text{align}}=0.08$ m on KITTI. The bottom-left area features a non-overlapping region. Colored circles visualize the neighborhood ranges of the corresponding scales.
(b) The case of poor registration with $E_{\text{align}}=2.46$ m on nuScenes. The red circle zooms in on a challenging region where larger neighborhoods are preferred.
@@
We report an ablation on the radius of the neighborhood and multiscale attention, and summarize the results in Table 2. Single-scale models prefer different radii per dataset: 7.5 m in nuScenes (high overlap) and 2.5 m in KITTI (low/moderate overlap), revealing the limitation of a single scale. Plain multiscale concatenation already improves over single-scale on nuScenes (noisy) and KITTI, but our point-wise scale-selector attention achieves the lowest RMSE on all datasets with only approximately 0.06% more parameters.
@@ Appendix A Feature Extraction
For each pair $(P^{\mathrm{S}},P^{\mathrm{R}})$, we first transform both point clouds into a common coordinate frame and apply farthest point sampling (FPS) [14] to each of them to select the same number of anchor points from $P^{\mathrm{S}}$ and $P^{\mathrm{R}}$, respectively.
Differential Entropy. Following CorAl [13], we use differential entropy as features. Let $\mathcal{N}_{P^{\rm S}}$ and $\mathcal{N}_{P^{\rm R}}$ denote the points from the source and reference scans within the neighborhood of anchor point $p$, respectively. Let the joint neighborhood be $\mathcal{N}_{\mathrm{joint}}=\mathcal{N}_{P^{\rm S}}\cup\mathcal{N}_{P^{\rm R}}$. We also use $\mathcal{N}_{\mathrm{sep}}$ to refer to $\mathcal{N}_{P^{\mathrm{S}}}$ if the anchor point belong to $P^{\mathrm{S}}$, and to $\mathcal{N}_{P^{\mathrm{R}}}$ if it belongs to $P^{\mathrm{R}}$. Assuming that each neighborhood has an approximately isotropic Gaussian distribution, its differential entropy has a closed form.
@@
where $\Sigma$ is the sample covariance of the points in neighborhood $\mathcal{N}$. Well-aligned point cloud pairs yield small gaps between $H(\mathcal{N}_{\mathrm{joint}})$ and $H(\mathcal{N}_{\mathrm{sep}})$, whereas misalignment inflates dispersion when the two sets are merged, increasing the gaps.
@@
$\qquad-\tfrac{1}{2}\Big[W_{\lambda}(\mathcal{N}_{P^{S}},\mathcal{N}_{P^{S}})+W_{\lambda}(\mathcal{N}_{P^{R}},\mathcal{N}_{P^{R}})\Big]~,$ (9)